Solve the ODE xex2+ydx=ydy


1
Expert's answer
2021-11-25T19:05:01-0500

"xe^{x^2+y}dx=dy\\
xe^{x^2}e^ydx=dy\\
xe^{x^2}dx=e^{-y}dy\\
\text{Integrating both side, we get}\\
put x^2=t\\
\implies 2xdx=dt\\
\frac{1}{2}\int e^tdt=\int e^{-y}dy\\
\frac{1}{2}e^t+C=-e^{-y}\\
\text{This is the required answer.}"